OpenStudy (anonymous):

Solve for p: 10p^3-5 ------- =0 p+1

OpenStudy (asnaseer):

is this your question?\[\frac{10p^3-5}{p+1}=0\]

OpenStudy (anonymous):

yes

OpenStudy (raden):

cross product the fraction, so we get : 10p^3-5 = 0(p+1) 10p^3-5 = 0 solve for x

OpenStudy (raden):

sorry, for p